OpenStudy (anonymous):

Profit Margin-Consumer Bus Math Could anyone explain how i could find this? Wes offered to pay $400 over the dealer's cost on a Knockout SX, which had a sticker price of $7,850. The dealer's profit margin is 10%. How much less would the dealer's profit be if she accepted Wes' offer?

jimthompson5910 (jim_thompson5910):

sticker price : $7,850 10% of this is 0.10*7850 = $785 so if the dealer got what she wanted (ie the 10% profit), then she would get 7850+785 = 8,635 dollars

jimthompson5910 (jim_thompson5910):

but...Wes only offers to pay $400 over the dealer cost (and not $785 over) so the dealer would lose 785 - 400 = 385 dollars if she took Wes' offer

OpenStudy (anonymous):

Oh ok i see now, at first i thought that it meant 400 more not 400 instead of 785, thank you so much

jimthompson5910 (jim_thompson5910):

sure thing